# BOPIM: Bayesian Optimization for influence maximization on temporal networks

### v2

@@ 3.3.1 Settings
We simulate the influence diffusion process and select seed sets for IM using $\mathsf{BOPIM}$. The main metrics of comparison are the number of influenced nodes at the conclusion of the spreading process, as well as the total algorithmic run time. We compare $\mathsf{BOPIM}$ with a greedy algorithm as this is considered the “gold standard” in IM. Erkol et al., (2022) proved that the SI model of diffusion is monotone and sub-modular so we use the CELF (Leskovec et al.,, 2007) step to increase the speed of the greedy algorithm. We also compare with random sampling seed nodes (Random), randomly sampling proportional to a nodes’ degree (Random Degree) and Dynamic Degree (Murata and Koga,, 2018). All experiments are conducted under the ex post assumption where the future evolution of the network is known when selecting seed nodes as in Murata and Koga, (2018); Osawa and Murata, (2015); Aggarwal et al., (2012); Erkol et al., (2020). While this assumption may be unrealistic in some cases, the main goal of this work is to propose a BO scheme for IM rather than tackling the ex ante problem (Yanchenko et al.,, 2023). We leave this as an important avenue of future work. Finally, to compare the different methods, we vary the number of seed nodes $(k)$, number of snapshots $(T)$, and infection parameter $(\lambda)$.
3.3.2 Results
@@
First, we fix $T=10$ and $\lambda=0.05,0.05,0.01,0.01$ for the Reality, Hospital, Bluetooth and Conference 2 datasets, respectively, while varying $k$. For this setting, we record both the influence spread and computing time with the results in Figures 2 and 3. For each data set, the proposed BO method yields comparable influence spread to the greedy algorithm. For Hospital and Conference 2, in particular, the spreads are almost indistinguishable between the two methods. $\mathsf{BOPIM}$ also yields a larger influence spread then Dynamic Degree in almost all settings. While Random Degree performs similarly to Dynamic Degree for some settings, Random always performs the worst. In terms of computing time, the proposed method has a large advantage over the greedy algorithm. In each case, the computational time greatly increases with $k$ for the greedy algorithm, whereas that of the proposed method increases more slowly. For large $k$, the proposed method is as much as five times faster than the greedy algorithm.
@@
Keeping the same values of $\lambda$ as before, we now fix $k$ such that roughly 5% of the total number of nodes are initially activated. We vary the number of aggregations of the temporal network, where small and large $T$ correspond to granular and fine aggregations, respectively. While the “real-time” that the network diffuses remains the same, we can vary the number of aggregations to see how this affects the influence spread. The results are in Figure 4. The trends are similar to that of varying $k$. Across networks and $T$, the proposed method yields comparable influence spread to that of the greedy algorithm, performing especially well in the Reality and Hospital networks. Dynamic Degree and Random Degree again yield consistently smaller influence spreads than both $\mathsf{BOPIM}$ and Greedy. In general, the total spread increases with $T$ as there are more rounds for influence to diffuse, but this trend need not be monotonic as different aggregations changes when certain nodes have edges. Indeed, in the Reality network, Dynamic Degree’s results are quite sensitive to $T$ as compared to the proposed method.
@@
Finally, we fix $k$ at 5% of $n$ and $T=10$ and vary the infection probability $\lambda$. The results are in Figure 5 and are similar for each network: as $\lambda$ increases, the total influence spread also increases. The distinction between different methods is least pronounced in this setting, but $\mathsf{BOPIM}$ has almost identical influence spreads to Greedy in all settings and networks.
@@ 3.4 Uncertainty quantification
The proposed method also provides measures of uncertainty for the optimal seed sets. Uncertainty quantification (UQ) allows for a richer understanding of the outputted solution and can answer questions such as: Are there many seed sets which yield near optimal spread? Can the selection of seed nodes be viewed on a spectrum as opposed to binary inclusion/exclusion? We take a first step towards answering these questions with the proposed method.
There are two main ways that we can use the $\mathsf{BOPIM}$ output for UQ. First, recall that we use GP regression to model the objective function. In (3), $\beta_{i}$ corresponds to the inclusion ($x_{i}=1$) or exclusion ($x_{i}=0$) of node $i$ in the seed set, and we obtain the complete posterior distribution for each coefficient. For the acquisition function step of $\mathsf{BOPIM}$, we only use the median of this posterior distribution. We can leverage the entire distribution, however, to obtain a clearer picture of a node’s importance.
@@
To do this, we save the final posterior distribution of $\bm{\beta}$ in $\mathsf{BOPIM}$ for the Reality network with $k=5$, $T=10$, $\lambda=0.05$ (15,000 MCMC samples, 5,000 burn-in). In Figure 6(a), we report the box-plot for the posterior distribution of $\beta_{i}$ for each node $i\in\{1,\dots,n\}$. The largest posterior distributions correspond with the ideal nodes for the seed set. From the figure, nodes 2, 3, 38 and 59 have noticeably larger values of the posterior distribution than that of the other nodes. Thus, while $k=5$ nodes are chosen for the seed set, it may be that nodes 2, 3, 38 and 59 are more influential than the other node included in the seed set.
@@
While $\mathsf{BOPIM}$ seeks the global optimum of our objective function, in practice, running the algorithm multiple times will give different optimal seed sets due to the randomness inherent in the algorithm. We can leverage this randomness for another way to quantify the uncertainty in the optimal seed sets. We now run the algorithm for 50 MC simulations (1,500 MCMC samples with 500 burn-in), and for each iteration, we find the optimal seed set. Then we can average these results to find the proportion of times that each node is assigned to the optimal seed set. These proportions are reported in Figure 6(b). For example, node 2 was selected in the seed set in 100% of the iterations, thus making it the most influential node. While nodes 1, 2 and 3 have noticeably larger proportions than all other nodes, the fourth and fifth largest (6 and 9) have much smaller proportions than that of those three.
@@
Finally, we note some differences between the two methods of UQ. Since Figure 6(b) is based on a single iteration, the results are subject to more stochastic variability, whereas much of this effect is averaged out in Figure 6(b). For example, nodes 38 and 59 appear to be key candidates for the seed set based on their large posterior distribution, but once we average over the results we see that these very rarely appears in the top $k$ of all nodes. On the other hand, the posterior mean of nodes 9 and 6 are close to zero in this one iteration, but they are the fourth and fifth most important nodes, respectively, when we re-run the algorithm many times. Regardless, these two figures together give us a good sense of the uncertainty in our optimal seed set and show that there are multiple sets of $k$ nodes which likely yield similar influence spreads.
@@
4 Discussion and conclusion
In this work, we apply a Bayesian Optimization framework to solve the influence maximization problem on temporal networks. The proposed $\mathsf{BOPIM}$ algorithm models the influence spread function via Gaussian Process regression and Horseshoe shrinkage prior. A simple mean function quantifies the importance of each nodes in the seed set, while our kernel function computes the correlation between seed sets based on their number of common neighbors. Experiments showed that despite being such a simple surrogate, the statistical model’s estimated influence spread is, on average, within a few nodes to the true spread. In the IM simulations, the proposed method yields influence spreads comparable to that of seed sets from a greedy algorithm. The BO approach, however, is much faster than the greedy algorithm. While in some settings, Dynamic Degree or Random Degree may have been similar to $\mathsf{BOPIM}$, the proposed algorithm consistently yielded the best results across all networks and settings. Moreover, the UQ found evidence that there are many seed sets which yield comparably optimal influence spread, thus implying that the objective function is relatively “flat.”
@@ 4 Discussion and conclusion
The posterior sampling scheme is another area where future work could be needed. In this work, we used a fully Gibbs approach to sample from the posterior distribution. As the size of the network increases, however, this step will become slower and more difficult. Thus, alternative approaches like Variational Bayes (Kingma and Welling,, 2013), where direct sampling from the posterior distribution is replaced with an optimization problem, may be necessary for computational advantages. Another area of extension is considering more complicated surrogate functions that take into account interactions between nodes in the seed set (Baptista and Poloczek,, 2018). This should lead to better performance that may even exceed the greedy algorithm. The number of parameters in the model, however, would be $O(n^{2})$ which could lead to computational challenges. Considering the ex ante IM task is another important open-problem.
Acknowledgements
This work was conducted while the author was on a Japan Society for the Promotion of Science (JSPS) fellowship. The author would also like to thank the editor and two anonymous reviewers whose comments greatly improved the quality of the manuscript.
@@ Acknowledgements
  1. 1.
    ${\bf S}=diag(\lambda_{1}^{2}\tau^{2},\dots,\lambda_{n}^{2}\tau^{2})$.
  2. 2.
    Sample $\sigma^{2}|\bm{\beta},\bm{\lambda},\bm{\nu},\tau^{2},,{\bf Y}\sim\mathsf{IG}(a_{1}+(N_{0}+n)/2,b_{1}+(({\bf Y}-{\bf X}\bm{\beta})^{T}{\bf K}^{-1}({\bf Y}-{\bf X}\bm{\beta})+\bm{\beta}^{T}{\bf S}^{-1}\bm{\beta})/2)$.
  3. 3.
    Sample $\lambda_{j}^{2}|\bm{\beta},\bm{\nu},\tau^{2},\sigma^{2}\sim\mathsf{IG}(1,\nu_{j}^{-1}+\beta^{2}_{j}/(2\tau^{2}\sigma^{2})$, $j=1,\dots,n$
  4. 4.
    Sample $\tau^{2}|\bm{\beta},\bm{\lambda},\sigma^{2},\xi\sim\mathsf{IG}((n+1)/2,\xi^{-1}+\sum_{k=1}^{n}\beta^{2}_{k}/(2\lambda^{2}_{k}\sigma^{2}))$.
  5. 5.
    Sample $\nu_{j}|\lambda_{k}\sim\mathsf{IG}(1,1+\lambda_{j}^{-2})$ for $j=1,\dots,n$.
  6. 6.
    Sample $\xi|\tau^{2}\sim\mathsf{IG}(1,1+\tau^{-2})$.
Greedy algorithm for $EI(\cdot)$
Result: Optimum seed set $\tilde{\bm{x}}$
Input: seed set size $k$, current optimum $\tilde{y}$, sampled sets ${\bf X}$, observations $\bm{y}$, ;
Initialize seed set $S$

### v4

@@ 3.2 Settings
We simulate the influence diffusion process and select seed sets for IM using $\mathsf{BOPIM}$. The main metrics of comparison are the proportion of nodes influenced at the conclusion of the spreading process, as well as the total algorithmic run time. We consider the proposed algorithm with both Hamming and Jaccard kernel, denoted $\mathsf{BOPIM-H}$ and $\mathsf{BOPIM-J}$, respectively, with $N_{0}=5$ and $B=20$. A greedy algorithm is also used, considered the “gold standard” in IM. Erkol et al., (2022) proved that the SI model of diffusion is monotone and sub-modular so we use the CELF (Leskovec et al.,, 2007) step to increase the speed of the greedy algorithm. We also compare with random sampling seed nodes (Random), randomly sampling proportional to a nodes’ degree (Random Degree) and Dynamic Degree (Murata and Koga,, 2018).
All experiments are conducted under the ex post assumption where the future evolution of the network is known when selecting seed nodes as in Murata and Koga, (2018); Osawa and Murata, (2015); Aggarwal et al., (2012); Erkol et al., (2020). While this assumption may be unrealistic in some cases, the main goal of this work is to propose a BO scheme for IM rather than tackling the ex ante problem (Yanchenko et al.,, 2023). We leave this as an important avenue of future work. Finally, to compare the different methods, we vary the number of seed nodes $(k)$, granularity of snapshots $(T)$, and infection parameter $(\lambda)$. Results are averaged over 25 Monte Carlo (MC) replications (save Greedy and Dynamic Degree), and the influence spread results are reported with error bars.
3.3 Results
First, we fix $T=10$ and $\lambda=0.05,0.05,0.01,0.01$ for the Reality, Hospital, Bluetooth and Conference 2 datasets, respectively, while varying $k$. In practice, $\lambda$ should be carefully chosen alongside domain experts, based on the particular scenario, but for illustrative purposes, we have chosen these values. For this setting, we record both the influence spread and computation time to select the seed set, with the results in Figures 2 and 3.
@@ 3.3 Results
For each data set, the proposed BO method yields comparable influence spread to the greedy algorithm, especially with the Hamming kernel. For Reality, Hospital and Conference 2, in particular, the spreads are almost indistinguishable between $\mathsf{BOPIM-H}$ and greedy. Both $\mathsf{BOPIM}$ methods yield similar influence spreads, except for the Reality network, and also exceed those of Dynamic Degree in all settings. While Random Degree performs similarly to Dynamic Degree for some settings, Random always performs the worst. In terms of computing time, the proposed method has a large advantage over the greedy algorithm. In each case, the computational time greatly increases with $k$ for the greedy algorithm, whereas that of the proposed method increases more slowly. For large $k$, the proposed method is as much as ten times faster than the greedy algorithm (Bluetooth network).
@@
Keeping the same values of $\lambda$ as before, we now fix $k$ such that roughly 5% of the total number of nodes are initially activated. We vary the number of aggregations of the temporal network, where small and large $T$ correspond to granular and fine aggregations, respectively. While the “real-time” that the network diffuses remains the same, we can vary the number of aggregations to see how this affects the influence spread. The results are in Figure 4.
The trends are similar to that of varying $k$. Across networks and $T$, the proposed method using Hamming kernel yields comparable influence spread to that of the greedy algorithm, again performing especially well in the Reality, Hospital and Conference 2 networks. We also see that the Hamming kernel noticeably outperforms the Jaccard kernel in the Reality, Hospital and Conference 2 networks. Dynamic Degree and Random Degree yield consistently smaller influence spreads than both $\mathsf{BOPIM-H}$ and Greedy. In general, the total spread increases with $T$ as there are more rounds for influence to diffuse, but this trend need not be monotonic as different aggregations changes when certain nodes have edges. Indeed, in the Reality network, Dynamic Degree’s results are quite sensitive to $T$ as compared to the proposed method.
@@
Finally, we fix $k$ at about 5% of $n$ and $T=10$ and vary the infection probability $\lambda$. The results are in Figure 5 and are similar for each network: as $\lambda$ increases, the total influence spread also increases. The distinction between different methods is least pronounced in this setting, but both versions of $\mathsf{BOPIM}$ have almost identical influence spreads to Greedy in all settings and networks.
@@ 4 Uncertainty quantification
One of the major advantages of the BO framework is that it naturally provides uncertainty quantification (UQ) results. Indeed, $\mathsf{BOPIM}$ is one of the first such methods for the IM problem. UQ allows for a richer understanding of the outputted solution and can answer questions such as: Are there many seed sets which yield near optimal spread? Can the selection of seed nodes be viewed on a spectrum as opposed to binary inclusion/exclusion? We take a first step towards answering these questions with the proposed method in this section by discussing two ways that $\mathsf{BOPIM}$ can be leveraged for UQ.
4.1 Posterior distribution of node contributions
4.1.1 Modified mean function
Recall that the mean function in Section 2.3.1 only included an intercept term. While this yielded seed sets with large influence spread, if we are interested in uncertainty quantification, it can be helpful to consider a more richly parameterized mean function.
Specifically, let the mean function $\mu(\cdot)$ be parameterized by $\bm{\beta}$. Motivated by Baptista and Poloczek, (2018), we assume that the surrogate model is linear in its parameters, i.e.,
$\mu(\bm{x}_{i})=\sum_{j=1}^{n}x_{ij}\beta_{j}=\bm{x}_{i}\bm{\beta}$ (7)
$\beta_{j}|\lambda_{j}^{2},\tau^{2},\sigma^{2}$ $\sim\mathsf{Normal}(0,\eta_{j}^{2}\tau^{2}\sigma^{2})$
$\eta_{j}^{2},\tau^{2}$ $\sim\mathsf{C}^{+}(0,1)$ (8)
where $j=1,\dots,n$ and $\mathsf{C}^{+}(0,1)$ is the standard half-Cauchy distribution. Thus, $\tau^{2}$ represents the global shrinkage and $\eta_{j}^{2}$ controls the local shrinkage of each coefficient. For the Gibbs sampler, we leverage an auxiliary variable formulation (Makalic and Schmidt,, 2015). Please see the Supplemental Materials for complete MCMC details. We also considered the R2D2 prior (Zhang et al.,, 2022; Yanchenko et al.,, 2025) and Dirichlet-Laplace (Bhattacharya et al.,, 2015), but the results were similar, so we focus on the HS prior.
4.1.2 Coefficient interpretation
4.1.3 Results
@@
We can leverage the posterior distribution of $\bm{\beta}$ to quantify the uncertainty of a node’s importance. To do this, we save the final posterior distribution of $\bm{\beta}$ in $\mathsf{BOPIM}$ for the Reality network with $k=5$, $T=10$, $\lambda=0.05$ (6,000 MCMC samples, 1,000 burn-in) and Hamming kernel. In Figure 6(a), we report the box-plot for the posterior distribution of $\beta_{i}$ for each node $i\in\{1,\dots,n\}$. The largest posterior distributions correspond with the ideal nodes for the seed set. From the figure, nodes 1, 2, and 3 have noticeably larger values of the posterior distribution than those of the other nodes. Thus, while $k=5$ nodes are chosen for the seed set, it may be that nodes 1, 2, and 3 are more influential than the other node included in the seed set.
4.2 Proportion of iterations selected for seed set
@@
While $\mathsf{BOPIM}$ seeks the global optimum of our objective function, in practice, running the algorithm multiple times will give different optimal seed sets due to the randomness inherent in the algorithm. We can leverage this randomness as another way to quantify the uncertainty in the optimal seed sets. We now run the algorithm for 50 MC simulations (2,500 MCMC samples with 500 burn-in), and for each iteration, we find the optimal seed set. Then we average these results to find the proportion of times that each node is assigned to the optimal seed set. Note that this measure of uncertainty holds regardless of the mean function, so we revert back to the intercept-only model for the following experiment. These proportions are reported in Figure 6(b). For example, node 2 was selected in the seed set in 100% of the iterations, thus making it the most influential node. While nodes 2 and 3 have noticeably larger proportions than all other nodes, the third through fifth largest (6, 8 and 9) have much smaller proportions than that of those two.
@@
We can note some differences between the two methods of UQ. Since Figure 6(a) is based on a single iteration, the results are subject to more stochastic variability, whereas much of this effect is averaged out in Figure 6(b). For example, node 1 has the second largest posterior distribution, but is only the sixth most likely node to appear in the seed set when averaged over several iterations. On the other hand, the posterior mean of nodes 9 and 6 are close to zero, but they are the third and fourth most important nodes, respectively, when we re-run the algorithm many times. Regardless, these two figures together give us a good sense of the uncertainty in our optimal seed set and show that there are multiple sets of $k$ nodes which likely yield similar influence spreads.
@@
5 Conclusion
In this work, we apply a Bayesian Optimization framework to solve the influence maximization problem on temporal networks. The proposed $\mathsf{BOPIM}$ algorithm models the influence spread function via Gaussian Process regression. In particular, we propose two kernel functions, one based on the Hamming distance and the other on the Jaccard coefficient. We also adopted the Expected Improvement acquisition function, optimizing this with a greedy algorithm which accounts for the cardinality constraints. Finally, we demonstrated $\mathsf{BOPIM}$’s utility in uncertainty quantification (UQ).
In all experiments, the proposed method yields influence spreads comparable to that of seed sets from a greedy algorithm. The BO approach, however, is much as ten times faster than the greedy algorithm. Additionally, we consistently found that the Hamming kernel was equivalent to, or outperformed, the Jaccard kernel. This result is somewhat surprising as we expected the Jaccard kernel to perform better since it explicitly accounts for the graph structure, unlike the Hamming distance. Additionally, the UQ results gave rich insights into the importance of nodes in the seed set. In particular, we found evidence that there are many seed sets which yield comparably optimal influence spread, thus implying that the objective function is relatively “flat.”
@@ 5 Conclusion
Another important area for future work is constructing the kernel function. Indeed, the surprising results of the Hamming distance’s superior performance indicates that the role of the kernel function in this problem is not fully understood. One potential direction comes from Automatic Relevance Detection (ARD) (Wipf and Nagarajan,, 2007; Neal,, 2012). This approach enforces sparsity in the model via the covariance function and has already enjoyed success in BO papers (e.g., Papenmeier et al.,, 2023). Finally, considering the ex ante IM task is another important open-problem.
@@
Supplemental methods and results
Proof of positive semi-definiteness of Hamming kernel (Section 1); discussion of COMBO algorithm’s applicability in IM (Section 2); greedy algorithm to optimize acquisition function (Section 3); relationship between greedy algorithm and Trust regions (Section 4); large network empirical results (Section 5); surrogate function validation results (Section 6); sampling scheme ablation study (Section 7); robustness experiments with $k$ and $T$ (Section 8); Horseshoe prior Gibbs sampler (Section 9)
Python code and data
Python code and data to replicate all results in Section 3 and re-create Figures 2 - 5. Code is also available on GitHub: https://github.com/eyanchenko/BOPIM.
Acknowledgments
This work was partially conducted while the author was on a Japan Society for the Promotion of Science (JSPS) fellowship. This work was also supported by JSPS KAKENHI grant 24K22613. The author would like to thank Tsuyoshi Murata and Petter Holme for their mentorship, as well as the editors and anonymous reviewers whose comments greatly improved the quality of the manuscript. Finally, ChatGPT-4o was used in the Hamming kernel proof to notice the connection with Gram matrices.
Disclosures
The author reports there are no competing interests to declare.
Supplemental Materials
This Supplemental Materials contains the following sections:
  1. 1.
    Proof of positive semi-definiteness of Hamming kernel
  2. 2.
    Further discussion of the COMBO algorithm in relation to the IM problem
  3. 3.
    The greedy algorithm to optimize the acquisition function
  4. 4.
    Discussion on the relationship between the greedy algorithm and Trust regions
  5. 5.
    Experiment comparing $\mathsf{BOPIM}$ with competing methods on a larger network
  6. 6.
    Surrogate function validation results
  7. 7.
    Ablation study on the sampling scheme for the initial observations in the BO algorithm
  8. 8.
    Robustness experiments when varying the training/testing values of $k$ and $T$
  9. 9.
@@ Acknowledgments
1. Proof of Hamming Kernel being positive semi-definite
First, we assume that $k<\frac{1}{2}n$. Recall that $\kappa(\bm{u},\bm{v})=1-\frac{1}{2k}d_{H}(\bm{u},\bm{v})$ for $i\neq j$, and $1+\delta$ if $i=j$. We want to prove that this function is positive semi-definite (psd) for all $\bm{u},\bm{v}\in\{0,1\}^{n}$ where $\sum_{i}u_{i}=\sum_{i}v_{i}=k$.
First, we can re-write the kernel as
$\kappa(\bm{u},\bm{v})=\frac{2k-n}{2k}+\frac{1}{2k}\sum_{i=1}^{n}\mathbb{I}(u_{i}=v_{i}),$
and note that the value is always positive due to the cardinality constraint on the input vectors. Then
$\sum_{i=1}^{n}\mathbb{I}(u_{i}=v_{i})=\sum_{i=1}^{n}\mathbb{I}(u_{i}=v_{i}=1)+\sum_{i=1}^{n}\mathbb{I}(u_{i}=v_{i}=0)=\bm{u}^{T}\bm{v}+({\bf 1}_{n}-\bm{u})^{T}({\bf 1}_{n}-\bm{v})$
where ${\bf 1}_{n}$ is the vector of one’s of length $n$. Thus, we can re-write the kernel as
$\kappa(\bm{u},\bm{v})$ $=\frac{2k-n}{2k}+\frac{1}{2k}\bm{u}^{T}\bm{v}+\frac{1}{2k}({\bf 1}_{n}-\bm{u})^{T}({\bf 1}_{n}-\bm{v})$
$=\frac{2k-n}{2k}+\frac{1}{2k}\bm{u}^{T}\bm{v}+\frac{n}{2k}-\frac{k}{2k}-\frac{k}{2k}+\frac{1}{2k}\bm{u}^{T}\bm{v}$
$=\frac{1}{k}\bm{u}^{T}\bm{v}.$
Thus, the covariance matrix is a scaled Gram matrix, implying positive $s$quare
2. COMBO algorithm
In this section, we briefly discuss $\mathsf{COMBO}$ and further explain why it is inapplicable to our IM problem. The main idea behind $\mathsf{COMBO}$ is to construct a kernel function between combinatorial inputs based on the combinatorial graph. [^4] To attempt to avoid confusion between the combinatorial graph and the original graph where the information diffusion occurs, we will refer to the former as being made up of vertices, and the latter as possessing nodes. In this graph, each vertex represents a possible (combinatorial) input to the objective function of interest, and edges connect vertices which differ only by a single input. In our context, this means that each vertex of the graph represents a possible seed set, and there is an edge between vertices if they differ by a single node in their respective seeds sets. To fix ideas, we construct the combinatorial graph for a toy example where we have $n=4$ nodes and a seed set of size $k=2$. Then the possible seed sets are $S\in\{1100,1010,1001,0110,0101,0011\}$ where the elements in $S$ have a 1 in position $i$ if node $i$ in included in the seed set, and 0 otherwise for $i=1,2,\dots,n=4$. For example, $1100$ means that nodes 1 and 2 are in the seed set. Then there is an edge between vertices if they share a seed node. The combinatorial graph can be seen Figure 7.
Figure 7: Combinatorial graph for possible seed sets with $n=4$ and $k=2$. Vertex label corresponds to the nodes assigned to the seed set where a 1 in position $i$ indicates that node $i$ is included in the seed set, and 0 otherwise. There is an edge between vertices which differ by only one node in the seed set.
Figure 8: In (a), we have the sub-graphs $\mathcal{G}_{i}$ corresponding to each node $i$. In (b), we take the graph Cartesian product of these sub-graphs to obtain the combinatorial graph, $\mathcal{G}_{C}$.
Now, to construct the combinatorial graph in general, the authors propose to use the graph Cartesian product. Specific to our problem, let $\mathcal{G}_{i}$ be the sub-graph corresponding to node $i$ for $i=1,\dots,n=4$. Then $\mathcal{G}_{i}$ is a sub-graph with two vertices, where one vertex corresponds to node $i$ being included in the seed set, and the other vertex corresponding to node $i$ being excluded from the seed set, with an edge connecting these two vertices. We construct such a sub-graph for each node $i$, and then the combinatorial graph, $\mathcal{G}_{C}$ is the graph Cartesian product, i.e., $\mathcal{G}_{C}=\square_{i}\mathcal{G}_{i}$. If $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ are two graphs with vertices $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$, respectively, then the graph Cartesian product is $\mathcal{H}=H_{1}\square H_{2}$ where $\mathcal{H}$ has vertices $\mathcal{V}_{1}\times\mathcal{V}_{2}$ and vertices $(h_{1},h_{2})$ and $(h_{1}^{\prime},h_{2}^{\prime})$ have an edge if and only if $h_{i}=h_{i}^{\prime}$ and $h_{j}$, $h_{j}^{\prime}$ have an edge in $\mathcal{H}_{j}$ for $i,j=1,2$.
In Figure 8(a), we show the sub-graphs $\mathcal{G}_{i}$ where $i=1,2,\dots,4$. The vertex on the left-hand side corresponds to node $i$ being included in the seed set, while the right-hand side vertex corresponds to the node not being included in the seed set. In Figure 8(b), we show the graph Cartesian product $\mathcal{G}=\square_{i=1}^{4}\mathcal{G}_{i}$. The difference between the graphs in Figure 7 and Figure 8(b) is clear. Not only are the number of vertices and edges different, but the graph constructed using the graph Cartesian product has several vertices which correspond to inadmissible seed sets, e.g., $0100$ only seeds one node while $1101$ seeds three nodes. The discrepancy is because the graph Cartesian product does not preserve our cardinality constraint.
While, technically, $\mathsf{COMBO}$ only requires constructing the combinatorial graph, all results in Oh et al., (2019) directly rely on constructing this graph using the graph Cartesian product. Since we cannot construct the graph in this way for our specific problem, we are unable to use $\mathsf{COMBO}$ for the IM task. We consider it an interesting avenue of future work to extend the ideas of $\mathsf{COMBO}$ to constrained input spaces.
3. Greedy algorithm for $AEI(\cdot)$
Result: Optimal seed set $\bm{x}^{*}$
Input: number of nodes $n$, seed set size $k$, acquisition function $AEI(\cdot)$
Initialize seed set $\bm{x}^{*}$ such that $\sum_{i=1}^{n}x_{i}^{*}=k$